# Improving Early Detection of Gravitational Waves from Binary Neutron Stars Using CNNs and FPGAs

### v2

@@ Abstract
I Introduction
In recent years, Astrophysics and adjacent sciences have been expanding at an astronomical pace. A significant driver of this growth is the detection of gravitational waves (GWs), which are ripples in spacetime generated by astronomical cataclysms. On September 2015, the Laser Interferometer Gravitational-Wave Observatory (LIGO) [1] and Virgo [2] collaborations confirmed the existence of GWs by detecting a binary black hole (BBH) merger [3]. Since then, over 90 confident events have been detected in the past three observation runs by LIGO-Virgo collaboration [4, 5, 6, 7]. Among these astronomical events, only two binary neutron star (BNS) mergers were detected [8].
As gravitational wave (GW) detectors become increasingly sensitive with each upgrade, the enhanced capabilities of second-generation detectors and the future third-generation detectors, such as the Einstein Telescope (ET) [9] or Cosmic Explorer [10], will allow for more frequent detection of binary neutron star (BNS) mergers and other exotic systems. Nonetheless, this increase in sensitivity comes at increase in computational complexity, as more sources will be detected. In the era of ET, it is estimated that $8\times 10^{4}\,$yr-1 BBH [11] and $7\times 10^{4}\,$yr-1 BNS [12] will be detected, which could lead to over 400 daily GW merger events, as well as the detection of other exotic sources [13]. Moreover, while BNS GW signals are detectable for seconds with current detectors, they will be present for hours in the ET era.
The current state-of-the-art for detecting modelled gravitational wave (GW) signals, known as matched-filtering—a technique based on cross-correlating models, or templates, with GW detector data [14, 15, 16, 17, 18]—has led the LIGO-Virgo-KAGRA collaboration to develop several low-latency matched-filtering pipelines for producing real-time GW alerts [19, 20, 21, 22, 23]. While these methods have been successful, matched-filtering techniques are known to be computationally intensive, posing challenges for future detectors and significantly contributing to increased pollution.
@@ I Introduction
GPUs, widely used for deploying ML algorithms, are well-known for being costly, consuming large amounts of energy, having short lifetimes and potentially introducing high latencies within data transfer. In a low-latency framework, these characteristics could hamper the early detection of GW signals. An interesting alternative is to embed ML algorithms in field-programmable gate arrays (FPGAs). FPGAs show promise by having longer lifetimes and being more energy-efficient than state-of-the-art hardware. Furthermore, high-end FPGAs are faster than CPUs and GPUs [24], and low-end FPGAs are more affordable than CPUs or GPUs in the same tier. We choose the latter to explore their applicability and limitations.
@@
In this work, to build more sustainable and environmentally-friendly alternatives, we explore the performance of FPGAs with respect to GPUs and CPUs to detect the early inspiral of BNS GW signals with ML-based algorithms  [^1] The code is accessible in GitHub: https://github.com/anaismartins/deploying-custom-nn-in-fpga. This paper is structured as follows: in Section II, we address the state of the art in early warning with ML-based methods and motivate the need to improve the precision at a limited computational cost; in Section III, we describe the input data; in section IV, we describe the hardware employed and ML architectures, as well as learning and quantization strategies (reduction from floating-point to fixed-point representation); in Section V, we present the main findings of this work, comparing our model with the state of the art in terms of accuracy and computational cost; finally, Section VI presents the conclusions of this research with avenues for future research.
II Related work
@@
The challenges in GW research require innovative solutions, and ML has emerged as a crucial tool for addressing them due to its adaptability and transversality. In the past few years, researchers have explored different ML applications to GW data analysis, such as detection of modelled [25, 26] and unmodelled GW signals [27, 28, 29, 30, 31, 32], non-transient burst noise characterization [33, 34] and synthetic data generation [35, 36, 37, 38], among others. Refer to [39] for a review.
@@
Traditional GW searches use the inspiral part of the waveform, as early efforts have shown that it is possible to detect a signal with only a fraction of the waveform model with matched-filtering techniques [19]. The first early warning ML-based algorithm implemented ResNet50 [40] training on spectrograms containing solely the inspiral part of the waveform [41]. However, as noted in [42], this added pre-processing step slows the inference by $\sim 0.5\,$s. To address this, [42] suggested using GW detector time series in design sensitivity, and employing 1-dimensional convolutional neural networks (CNNs), eliminating the need for this pre-processing step and enhancing efficiency. This proof of concept work led to subsequent efforts in this line of research [43, 25, 44, 26, 45, 46].
@@
In particular, [26], building on the previous work of [42] and [25], extended this method to the simulated noise of third and fourth observing runs (O3 and O4), as well as real O3 data. While [26] is the current state of the art, some of the limitations of this study are described below:
@@
To overcome the limitations of the previous work, hereafter referred to as FindCNN, we implemented an algorithm inspired by WaveNet [47] named GWaveNet. Similarly to [48] and [30], this choice was motivated by the similitude between GW time series data and audio or speech signals. For a fair comparison, we reproduced FindCNN and trained and tested both models on the same simulated GW detector data.
FPGAs have been used for ML applications [49], also in high-energy physics [50, 51]. We highlight the pioneering work in [52] for the implementation of recurrent neural networks for anomaly detection in GW. As added value, in this work we evaluate the performance of the FindCNN and GWaveNet models in CPU, GPU and FPGA devices, as well as their time, energy consumption and economical cost.
III Data set and pre-processing
@@
In a real scenario, the negative class is much larger than the positive class, as only $\sim 100$ GW signals have been discovered at the present date. Thus, it is required to simulate GW signals, adding them to detector noise. This process is commonly referred to as “injections”, and they will be the positive class of this study. Hence, we construct a balanced data set of samples with only detector noise (negative class), and injections (positive class). Simulated Gaussian noise is generated with the average power spectral density of the third observing run (O3) for Hanford, Livingston and Virgo [^2] We use the PyCBC package [53]. In particular, the power spectral densities aLIGOaLIGO140MpcT1800545 and aLIGOAdVO3LowT1800545, an average of O3, respectively for the Advanced LIGO and Advanced Virgo detectors [54]. Refer to https://gwosc.org/O3/o3_details/ for technical details of the observing run..
As proposed in [26], to encompass all potential BNS systems, the component masses are uniformly distributed $\in[1,3]\,\text{M}_{\odot}$, and the waveform approximant SpinTaylorT4 [55] (used to model inspirals) was chosen. Moreover, the sources are uniformly distributed over the sky, including spin effects.
@@
As GWs have a weak amplitude, they are hard to distinguish from detector noise. For illustration, in Fig. 1 (top panel) we show the GW inspiral of a signal similar to GW170817 [8] in raw Advanced LIGO detector noise. To highlight its amplitude a common practice is to filter the data between $10$ and 100100100\,100Hz, and whiten it, i.e. make the signal Gaussian-like with uniform variance by removing all the correlation of the noise [56]. Afterwards, we normalize the data $\in[-1,1]$. As we can observe in Fig. 1, the inspiral phase lasts several minutes, so, as in [26], we pre-select 300300300\,300s of data. In the bottom panel, we show the relation between time and frequency of the source. Because of this relation, the difficulty of the input is governed by the maximum frequency $f_{max}$ within a time window. Samples with higher (lower) $f_{max}$ will be closer (further) to the merger, meaning that they will be detected later (earlier).
@@ IV Methodology
IV-A Hardware specifications
The manufacturing and use of state-of-the-art hardware are causing substantial damage to the environment and this is only predicted to increase, with the manufacturing of computer hardware predicted to take up more than 20% of global energy usage by 2030 [57] and computation energy predicted to hit the world’s energy production capacity by 2040 [58]. In this context, it is of interest to understand the computational cost of low-latency algorithms of large physics experiments, such as Advanced LIGO-Virgo detectors, as well as their lifetime under heavy usage. In the following, we provide the specifications of the hardware employed in this work.
    CPU: We use the AMD EPYC 7551P 32-Core Processor [59], which has an expected lifetime of around 5 to 10 years. We used the full 32 cores for testing.
    GPU: We used the NVIDIA Tesla V100 [60] for training and validation, and NVIDIA GeForce GTX 108 [61] for testing and energy consumption calculation. Their expected lifetime is around 3 to 5 years.
    FPGA: We use the AMD Kria KV260 Vision AI Starter Kit [62]. Its expected lifetime is around 10 to 15 years. We use the most recent and fastest pre-made architecture for this system-on-chip, B4096, with 4096 peak operations per cycle. Furthermore, we use 10 threads for testing, which we found to be the optimal amount.
Figure 2: Overview of the modified WaveNet convolutional architecture for a single data stream, adapted from [63]. We present $n=7$ 1-dimensional WaveNet modules composed of dilated causal convolutions (blue) with the number of filters $f$, kernel size $k$, dilation $d$ and padding $p$, batch normalization layers (orange), $\tanh$ and sigmoid $\sigma$ activations (green). Furthermore, we show $1\times 1$ convolutions (blue), dense layers (pink), and ReLU and Softmax activations (green).
@@ IV-A Hardware specifications
IV-B Basic components
IV-B1 Common details
@@
We implemented FindCNN and GWaveNet using the PyTorch package [64]. Both networks minimize the binary cross-entropy (BCE) loss in PyTorch, the BCEWithLogitsLoss i.e., Sigmoid activation integrated with BCE loss [65]. We decided to employ this loss instead of BCELoss since it is more numerically stable. Note that the original FindCNN [26] uses BCELoss instead.
IV-B2 FindCNN
@@
It is a 1-dimensional CNN [26] that starts with a batch normalization layer [66], to stabilize and accelerate the training process, followed by 5 convolutional blocks and two dense layers. The convolutional blocks are composed of a convolutional layer with stride 4, a ReLU activation function [67], and a MaxPooling layer [68] with stride 1. Each convolutional layer has successively kernel sizes $k$ $[16,8,4,8,16]$, and number of filters $f$ $[32,64,128,256,256]$. This structure allows capturing high-level features at earlier layers while capturing finer details at intermediate layers. Furthermore, the combination of large kernel $k$ and large filters allowed the last layers to capture high-level context with great expressiveness. Despite the compression of the input due to the usage of large kernel sizes $k$, the first dense layer is one of the most expensive ones, as it further compresses an input of $37,632$ data points to only $128$. As expected, this model is highly complex, having $6,179,303$ trainable parameters, mainly due to the length of the input and large dense layers.
IV-B3 GWaveNet
@@
WaveNet is an expressive CNN designed for the generation of high-fidelity speech audio [63]. The architecture is capable of handling long-range temporal dependencies, a characteristic desirable for GW data analysis, and in particular for BNS searches, due to their long inspirals. WaveNet-like models have been previously implemented in the field of GW searches [48, 30]. These implementations modify WaveNet for binary classification. This is a common point in the current investigation. However, previous WaveNet models were constructed as an ensemble with non-causal dilated convolutions, while we encode the information of all detectors in a single model with causal dilated convolutions.
@@
Causal dilated convolutions combine the power of dilation [63], which allows the NN to capture a larger context without increasing the computational load significantly, and causal convolutions, to effectively learn the time ordering of the data. Our experiments shown an improved performance when obeying the arrow of time with causal convolutions. These convolutions are encapsulated in WaveNet modules.
@@
We present the full architecture of GWaveNet, including the WaveNet modules in Fig. 2. We can observe how an input $x$ gets duplicated and each copy gets fed into a causal convolution block (blue), followed by a batch normalization layer (orange). Then, one output passes a $\tanh$ activation function (green), while the other passes a Sigmoid $\sigma$ activation function (green). This is known as gated activation [69], designed for capturing complex relations in sequential data. Afterwards, the two outputs are multiplied (grey), generating the output of the module, $y$. At this point, the initial input $x$ is added to the output $y$ to form the residual input $y_{res}$ that passes to the next module. On the other hand, the output $y$ is sent towards the final output of the network that will feed the BCEWithLogitsLoss. Before this step, all the $y$ outputs of the modules are summed, passing a $1\times 1$ convolution (blue) and two dense layers (pink) with ReLU activation functions (green) [67]. It is relevant to note that original WaveNet modules used $1\times 1$ convolutions for further compression, but due to the fixed number of filters $f$, this was not needed.
@@
IV-C Learning strategy
As we mentioned in Sections II and III, [26] implemented a curriculum learning strategy going down in maximum frequency $f_{max}$. For this aim, five data sets with different average $f_{max}$ are built (see Table I). In this way, the model first learns easier examples, closer to the merger, and slowly transitions to harder examples, further from the merger. Each data set is composed of $22,600$ samples for both classes, such that for each curriculum learning step we use $71\%$ for training, $9\%$ for validation and $20\%$ for testing.
In GW data analysis, it is common to measure the loudness of the signal in terms of signal-to-noise ratio (SNR) [18]. However, [25] and [26] introduced the concept of partial inspiral signal-to-noise ratio (PISNR) to describe the corresponding SNR of the partial waveform, seen by the CNN. Thus, we use this magnitude for comparison (Table I).

### v3

@@ Abstract
  • December 19, 2024
1 Introduction
In recent years, Astrophysics and adjacent sciences have been expanding at an astronomical pace. A significant driver of this growth is the detection of gravitational waves (GWs), which are ripples in spacetime generated by astronomical cataclysms. On September 2015, the Laser Interferometer Gravitational-Wave Observatory (LIGO) Aasi et al. (2015) and Virgo Acernese et al. (2015) collaborations confirmed the existence of GWs by detecting a binary black hole (BBH) merger Abbott et al. (2016). Since then, over 90 confident events have been detected in the past three observation runs by LIGO-Virgo collaboration Abbott et al. (2019, 2021, 2024, 2023a). Among these astronomical events, only two binary neutron star (BNS) mergers were detected Abbott et al. (2017).
As gravitational wave (GW) detectors become increasingly sensitive with each upgrade, the enhanced capabilities of second-generation detectors and the future third-generation detectors, such as the Einstein Telescope (ET) Punturo et al. (2010) or Cosmic Explorer Reitze et al. (2019), will allow for more frequent detection of binary neutron star (BNS) mergers and other exotic systems. Nonetheless, this increase in sensitivity comes at increase in computational complexity, as more sources will be detected. In the era of ET, it is estimated that $8\times 10^{4}\,$yr-1 BBH Iacovelli et al. (2022) and $7\times 10^{4}\,$yr-1 BNS Kalogera et al. (2021) will be detected, which could lead to over 400 daily GW merger events, as well as the detection of other exotic sources Baiotti et al. (2005). Moreover, while BNS GW signals are detectable for seconds with current detectors, they will be present for hours in the ET era.
The current state-of-the-art for detecting modelled gravitational wave (GW) signals, known as matched-filtering—a technique based on cross-correlating models, or templates, with GW detector data Messick et al. (2017); Sachdev et al. (2019); Tsukada et al. (2023); Usman et al. (2016); Allen et al. (2012)—has led the LIGO-Virgo-KAGRA collaboration to develop several low-latency matched-filtering pipelines for producing real-time GW alerts Sachdev et al. (2020); Nitz et al. (2020); Magee et al. (2021); Kovalam et al. (2022); Chaudhary et al. (2024). While these methods have been successful, matched-filtering techniques are known to be computationally intensive, posing challenges for future detectors and significantly contributing to increased CO2 pollution.
@@ 1 Introduction
GPUs, widely used for deploying ML algorithms, are well-known for being costly, consuming large amounts of energy, having short lifetimes and potentially introducing high latencies within data transfer. In a low-latency framework, these characteristics could hamper the early detection of GW signals. An interesting alternative is to embed ML algorithms in field-programmable gate arrays (FPGAs). FPGAs show promise by having longer lifetimes and being more energy-efficient than state-of-the-art hardware. Furthermore, high-end FPGAs are faster than CPUs and GPUs Liu et al. (2012), and low-end FPGAs are more affordable than CPUs or GPUs in the same tier. We choose the latter to explore their applicability and limitations.
@@
In this work, to build more sustainable and environmentally-friendly alternatives, we explore the performance of FPGAs with respect to GPUs and CPUs to detect the early inspiral of BNS GW signals with ML-based algorithms  [^3] The code is accessible in GitHub: https://github.com/anaismartins/deploying-custom-nn-in-fpga. This paper is structured as follows: in Section 2, we address the state of the art in early warning with ML-based methods and motivate the need to improve the precision at a limited computational cost; in Section 3, we describe the input data; in section 4, we describe the hardware employed and ML architectures, as well as learning and quantization strategies (reduction from floating-point to fixed-point representation); in Section 5, we present the main findings of this work, comparing our model with the state of the art in terms of accuracy and computational cost; finally, Section 6 presents the conclusions of this research with avenues for future research.
2 Related work
@@
The challenges in GW research require innovative solutions, and ML has emerged as a crucial tool for addressing them due to its adaptability and transversality. In the past few years, researchers have explored different ML applications to GW data analysis, such as detection of modelled Baltus et al. (2021b, 2022) and unmodelled GW signals Cavaglia et al. (2020); López Portilla et al. (2021); López et al. (2021); Meijer et al. (2024); Boudart and Fays (2022); Boudart (2023), non-transient burst noise characterization Bahaadini et al. (2018); Laguarta et al. (2023) and synthetic data generation Lopez et al. (2022b, a); Dooney et al. (2022, 2024), among others. Refer to Cuoco et al. (2021) for a review.
@@
Traditional GW searches use the inspiral part of the waveform, as early efforts have shown that it is possible to detect a signal with only a fraction of the waveform model with matched-filtering techniques Sachdev et al. (2020). The first early warning ML-based algorithm implemented ResNet50 He et al. (2015) training on spectrograms containing solely the inspiral part of the waveform Wei et al. (2021a). However, as noted in Baltus et al. (2021a), this added pre-processing step slows the inference by $\sim 0.5\,$s. To address this, Baltus et al. (2021a) suggested using GW detector time series in design sensitivity, and employing 1-dimensional convolutional neural networks (CNNs), eliminating the need for this pre-processing step and enhancing efficiency. This proof of concept work led to subsequent efforts in this line of research Wei et al. (2021b); Baltus et al. (2021b); Yu et al. (2021); Baltus et al. (2022); Reem et al. (2024); van Straalen et al. (2024).
@@
In particular, Baltus et al. (2022), building on the previous work of Baltus et al. (2021a) and Baltus et al. (2021b), extended this method to the simulated noise of third and fourth observing runs (O3 and O4), as well as real O3 data. While Baltus et al. (2022) is the current state of the art, some of the limitations of this study are described below:
@@
To overcome the limitations of the previous work, hereafter referred to as FindCNN, we implemented an algorithm inspired by WaveNet van den Oord et al. (2016b) named GWaveNet. Similarly to Wei et al. (2021c) and Meijer et al. (2024), this choice was motivated by the similitude between GW time series data and audio or speech signals. For a fair comparison, we reproduced FindCNN and trained and tested both models on the same simulated GW detector data.
FPGAs have been used for ML applications Sang et al. (2016), also in high-energy physics Khoda et al. (2022); Que et al. (2024). We highlight the pioneering work in Que et al. (2021) for the implementation of recurrent neural networks for anomaly detection in GW. As added value, in this work we evaluate the performance of the FindCNN and GWaveNet models in CPU, GPU and FPGA devices, as well as their time, energy consumption and economical cost.
3 Data set and pre-processing
@@
In a real scenario, the negative class is much larger than the positive class, as only $\sim 100$ GW signals have been discovered at the present date. Thus, it is required to simulate GW signals, adding them to detector noise. This process is commonly referred to as “injections”, and they will be the positive class of this study. Hence, we construct a balanced data set of samples with only detector noise (negative class), and injections (positive class). Simulated Gaussian noise is generated with the average power spectral density of the third observing run (O3) for Hanford, Livingston and Virgo [^4] We use the PyCBC package Biwer et al. (2019). In particular, the power spectral densities aLIGOaLIGO140MpcT1800545 and aLIGOAdVO3LowT1800545, an average of O3, respectively for the Advanced LIGO and Advanced Virgo detectors Abbott et al. (2023b). Refer to https://gwosc.org/O3/o3_details/ for technical details of the observing run..
As proposed in Baltus et al. (2022), to encompass all potential BNS systems, the component masses are uniformly distributed $\in[1,3]\,\text{M}_{\odot}$, and the waveform approximant SpinTaylorT4 Buonanno et al. (2003) (used to model inspirals) was chosen. Moreover, the sources are uniformly distributed over the sky, including spin effects.
@@
As GWs have a weak amplitude, they are hard to distinguish from detector noise. For illustration, in Fig. 1 (top panel) we show the GW inspiral of a signal similar to GW170817 Abbott et al. (2017) in raw Advanced LIGO detector noise. To highlight its amplitude a common practice is to filter the data between $10$ and 100100100\,100Hz, and whiten it, i.e. make the signal Gaussian-like with uniform variance by removing all the correlation of the noise Akhshi et al. (2020). Afterwards, we normalize the data $\in[-1,1]$. As we can observe in Fig. 1, the inspiral phase lasts several minutes, so, as in Baltus et al. (2022), we pre-select 300300300\,300s of data. In the bottom panel, we show the relation between time and frequency of the source. Because of this relation, the difficulty of the input is governed by the maximum frequency $f_{max}$ within a time window. Samples with higher (lower) $f_{max}$ will be closer (further) to the merger, meaning that they will be detected later (earlier).
@@ 4 Methodology
4.1 Hardware specifications
The manufacturing and use of state-of-the-art hardware are causing substantial damage to the environment and this is only predicted to increase, with the manufacturing of computer hardware predicted to take up more than 20% of global energy usage by 2030 News (2022) and computation energy predicted to hit the world’s energy production capacity by 2040 Office (2022). In this context, it is of interest to understand the computational cost of low-latency algorithms of large physics experiments, such as Advanced LIGO-Virgo detectors, as well as their lifetime under heavy usage. In the following, we provide the specifications of the hardware employed in this work.
    CPU: We use the AMD EPYC 7551P 32-Core Processor TechPowerUp (2024), which has an expected lifetime of around 5 to 10 years. We used the full 32 cores for testing.
    GPU: We used the NVIDIA Tesla V100 Microway, Inc. (2018) for training and validation, and NVIDIA GeForce GTX 1080 NVIDIA (2024) for testing and energy consumption calculation. Their expected lifetime is around 3 to 5 years.
    FPGA: We use the AMD Kria KV260 Vision AI Starter Kit Advanced Micro Devices, Inc. (Accessed 27 March 2024). Its expected lifetime is around 10 to 15 years. We use the most recent and fastest pre-made architecture for this system-on-chip, B4096, with 4096 peak operations per cycle. Furthermore, we use 10 threads for testing, which we found to be the optimal amount.
Figure 2: Overview of the modified WaveNet convolutional architecture for a single data stream, adapted from van den Oord et al. (2016c). We present $n=7$ 1-dimensional WaveNet modules composed of dilated causal convolutions (blue) with the number of filters $f$, kernel size $k$, dilation $d$ and padding $p$, batch normalization layers (orange), $\tanh$ and sigmoid $\sigma$ activations (green). Furthermore, we show $1\times 1$ convolutions (blue), dense layers (pink), and ReLU and Softmax activations (green).
@@ 4.1 Hardware specifications
4.2 Basic components
4.2.1 Common details
@@
We implemented FindCNN and GWaveNet using the PyTorch package A. et al. (2019). Both networks minimize the binary cross-entropy (BCE) loss in PyTorch, the BCEWithLogitsLoss i.e., Sigmoid activation integrated with BCE loss PyTorch (Accessed 24 April 2024). We decided to employ this loss instead of BCELoss since it is more numerically stable. Note that the original FindCNN Baltus et al. (2022) uses BCELoss instead.
4.2.2 FindCNN
@@
It is a 1-dimensional CNN Baltus et al. (2022) that starts with a batch normalization layer Ioffe et al. (2015), to stabilize and accelerate the training process, followed by 5 convolutional blocks and two dense layers. The convolutional blocks are composed of a convolutional layer with stride 4, a ReLU activation function Nair et al. (2010), and a MaxPooling layer PyTorch (Accessed 22 April 2024) with stride 1. Each convolutional layer has successively kernel sizes $k$ $[16,8,4,8,16]$, and number of filters $f$ $[32,64,128,256,256]$. This structure allows capturing high-level features at earlier layers while capturing finer details at intermediate layers. Furthermore, the combination of large kernel $k$ and large filters allowed the last layers to capture high-level context with great expressiveness. Despite the compression of the input due to the usage of large kernel sizes $k$, the first dense layer is one of the most expensive ones, as it further compresses an input of $37,632$ data points to only $128$. This model has $6,179,303$ trainable parameters, mainly due to the length of the input and large dense layers.
4.2.3 GWaveNet
@@
WaveNet is an expressive CNN designed for the generation of high-fidelity speech audio van den Oord et al. (2016c). The architecture is capable of handling long-range temporal dependencies, a characteristic desirable for GW data analysis, and in particular for BNS searches, due to their long inspirals. WaveNet-like models have been previously implemented in the field of GW searches Wei et al. (2021c); Meijer et al. (2024). These implementations modify WaveNet for binary classification. This is a common point in the current investigation. However, previous WaveNet models were constructed as an ensemble with non-causal dilated convolutions, while we encode the information of all detectors in a single model with causal dilated convolutions.
@@
Causal dilated convolutions combine the power of dilation van den Oord et al. (2016c), which allows the NN to capture a larger context without increasing the computational load significantly, and causal convolutions, to effectively learn the time ordering of the data. Our experiments shown an improved performance when obeying the arrow of time with causal convolutions. These convolutions are encapsulated in WaveNet modules.
@@
We present the full architecture of GWaveNet, including the WaveNet modules in Fig. 2. We can observe how an input $x$ gets duplicated and each copy gets fed into a causal convolution block (blue), followed by a batch normalization layer (orange). Then, one output passes a $\tanh$ activation function (green), while the other passes a Sigmoid $\sigma$ activation function (green). This is known as gated activation van den Oord et al. (2016a), designed for capturing complex relations in sequential data. Afterwards, the two outputs are multiplied (grey), generating the output of the module, $y$. At this point, the initial input $x$ is added to the output $y$ to form the residual input $y_{res}$ that passes to the next module. On the other hand, the output $y$ is sent towards the final output of the network that will feed the BCEWithLogitsLoss. Before this step, all the $y$ outputs of the modules are summed, passing a $1\times 1$ convolution (blue) and two dense layers (pink) with ReLU activation functions (green) Nair et al. (2010). It is relevant to note that original WaveNet modules used $1\times 1$ convolutions for further compression, but due to the fixed number of filters $f$, this was not needed.
@@
4.3 Learning strategy
As we mentioned in Sections 2 and 3, Baltus et al. (2022) implemented a curriculum learning strategy going down in maximum frequency $f_{max}$. For this aim, five data sets with different average $f_{max}$ are built (see Table 1). In this way, the model first learns easier examples, closer to the merger, and slowly transitions to harder examples, further from the merger. Each data set is composed of $22,600$ samples for both classes, such that for each curriculum learning step we use $71\%$ for training, $9\%$ for validation and $20\%$ for testing.
In GW data analysis, it is common to measure the loudness of the signal in terms of signal-to-noise ratio (SNR) Allen et al. (2012). However, Baltus et al. (2021b) and Baltus et al. (2022) introduced the concept of partial inspiral signal-to-noise ratio (PISNR) to describe the corresponding SNR of the partial waveform, seen by the CNN. Thus, we use this magnitude for comparison (Table 1).